### v1

# Maximal non-compactness of embeddings between sequence spaces

Abstract.
We will focus on studying the ball measure of non-compactness $\alpha(T)$ for various particular instances of embedding operators in sequence spaces. Our first main goal is to find necessary and sufficient conditions for an identity operator to be maximally non-compact. Next, we will focus on studying Lorentz sequence spaces $\ell^{p,q}$ and their basic properties. We will characterize the inclusions between Lorentz sequence spaces depending on the values of $p$ and $q$. Then we will try to determine exact values of the norms of the identity operators between these embedded spaces. Lastly, we will determine whether these identity operators are maximally non-compact by using our general theorems.
Key words and phrases:
embedding, maximal non-compactness, sequence spaces, normed linear spaces, Lorentz spaces
Introduction
Investigation of compactness of mappings is arguably one of the most important subdisciplines of analysis of all times, and its results have proved useful in several fields of mathematics including contemporary functional analysis. It is interesting not only from the theoretical point of view but, at the same time, it has a wide array of applications. However, especially in infinite-dimensional spaces, the mapping of interest is often not compact.
The difference between compactness and non-compactness of an operator is often too rough, and, therefore, it is worthwhile to study its various refinements. One particular example of such a refinement is offered by studying the so-called measure of non-compactness.
The measure of non-compactness, $\alpha(T)$, of a mapping $T$ acting from one quasinormed space to another, can attain any value between zero and the operator norm $T||$. The mapping $T$ is compact if and only if $\alpha(T)=0$, at least on complete spaces. On the other endpoint of the segment of possible values of $\alpha(T)$, that is, in the case when $\alpha(T)$ coincides with the operator norm, the mapping is called maximally non-compact, marking the worst possible case.
The concept of measure of non-compactness is thus a good device for quantifying how bad the non-compactness of a mapping is. The origins of the idea go back to the 1930 paper by Kuratovskii [11]. Since its introduction, it has been proven very useful, and found many quite diverse applications, see, for instance, [1, 6, 16]. The concept of measure of noncompactness is intimately connected with the limiting behavior of entropy numbers, and the $s$-numbers in general [5, 4]. The advances of the field led eventually to development of a formidable theory which found its way to important classical monographs such as [7, 8, 15, 17].
An important particular example of an operator whose compactness or non-compactness is studied is the identity. In such cases, we talk about embeddings of the corresponding spaces or classes. Measure of noncompactness of embeddings has been studied extensively, for example in connection of Sobolev embeddings and their applications to studying properties of solution to elliptic partial differential equation. The fact that here is an immense interest in studying these problems can be illustrated by the vast amount of papers including some very recent ones. To name just very few, see e.g. [3, 9, 14, 12, 13].
In sharp contrast to the extensive existing literature on compactness-related properties of function spaces stands the fact that almost nothing is known about analogous problems concerning sequence spaces. This paper is intended as a contribution towards filling this gap.
Our aim is to study questions about maximal non-compactness of embeddings of various sequence spaces. After stating several general statements, we shall focus on the particular scale of Lorentz sequence spaces.
The paper is structured as follows. In the next section we collect all relevant definitions and fix notation. In Section 2 we state and prove our first array of main results. We focus here on embeddings between general sequence spaces which are endowed at least with a quasi-norm. More precisely, we offer here a sufficient condition for the maximal non-compactness of an embedding between two sequence spaces included into $c_{0}$, an analogous result concerning rearrangement-invariant lattices, and also a sufficient condition for such an embedding to avoid maximal non-compactness. In our approach to the latter problem we employ the recent concept of span of an embedding.
We then turn our attention to Lorentz sequence spaces, which are collections of sequences determined by the two-parameter scale of Lorentz functionals. It is known that the Lorentz functional is always a quasinorm, which in very special cases turns to a norm or is equivalent to a norm. Lorentz spaces constitute a very interesting collection of sequence spaces, even slightly richer than their companions containing functions (note that, for example, spaces $\ell^{\infty,q}$ with finite $q$ are nontrivial).
In order to enable application of our general results to Lorentz spaces, we begin by showing that, for an arbitrary combination of $p,q\in[1,\infty]$ such that $\min\{p,q\}<\infty$, the space $\ell^{p,q}$ is a subset of $c_{0}$ by some elementary means. Next, we concentrate on characterizing the inclusions $\ell^{p_{1},q_{1}}\hookrightarrow\ell^{p_{2},q_{2}}$ between the Lorentz sequence spaces depending on the values of $p_{1},p_{2}$ and $q_{1},q_{2}$. Here, our point of departure are inequalities and inclusions mentioned in [10], which we extend and fill in details to complete the proofs and give a comprehensive picture. In most cases, we furthermore determine the exact values of the norms of the identity operators between these spaces. This knowledge is very useful in our last set of results, in which we determine whether embedding relations between Lorentz sequence spaces are maximally non-compact.
1. Preliminaries
Definition 1.1 (Ball measure of non-compactness).
Let $X$, $Y$ be (alpha/quasi)-normed linear spaces over $\mathbb{K}$ and let $T\colon X\rightarrow Y$ be a bounded mapping defined on $X$ and taking values in $Y$. The ball measure of non-compactness $\alpha(T)$ of $T$ is defined as the infimum of radii $r>0$ for which there exists a finite set of balls in $Y$ of radii $r$ that covers $T(B_{X})$, where $B_{X}$ is the closed unit ball on $X$. In other words,
$\alpha(T)=\inf\{r>0:T(B_{X})\subset\bigcup_{i=1}^{m}(y_{i}+rB_{Y})\;\;y_{i}\in Y\;\&\;m\in\mathbb{N}\}.$
Let us recall that given $X$ and $Y$ (alpha/quasi)-normed linear spaces over $\mathbb{K}$ and $T\colon X\rightarrow Y$ a bounded mapping defined on $X$ and taking values in $Y$ we define the operator norm $T||$ of $T$ as follows:
$\|T\|=\sup_{x\in B_{X}}\|T(x)\|.$
We will now formulate an elementary, but important inequality estimating the Ball measure of non-compactness.
Proposition 1.2.
Let $X$ and $Y$ be (alpha/quasi)-normed linear spaces over $\mathbb{K}$ and let $T\colon X\rightarrow Y$ be a bounded linear mapping defined on $X$ and taking values in $Y$. Then:
$0\leq\alpha(T)\leq\|T\|.$
Proof.
We will prove each of the inequalities separately:
We can see that $\alpha(T)\geq 0$ easily from the definition of the ball measure of non-compactness, because infimum of positive radii $r$ is always greater or equal to zero.
Take $\varrho>\|T\|$ arbitrarily. Then for every $x\in B_{X}$ one has $\|Tx\|\leq\|T\|<\varrho$, which in turn implies that $Tx\in B_{Y}(0,\varrho)$. Since $x$ was arbitrary, this implies that
$T(B_{X})\subset\bigcup_{i=1}^{m=1}(0+\varrho B_{Y}).$
The last inclusion is true for all $\varrho>\|T\|$. Consequently, $\alpha(T)\leq\|T\|$. ∎
Definition 1.3 (Maximal non-compactness).
Let $X$ and $Y$ be (alpha/quasi)-normed linear spaces over $\mathbb{K}$ and let $T\colon X\rightarrow Y$ be a bounded mapping defined on $X$ and taking values in $Y$. We say that $T$ is maximally non-compact if
$\alpha(T)=\|T\|.$
The following statement gives us the relationship between the value of the Ball measure of non-compactness of an operator and compactness itself.
Proposition 1.4.
Let $X$ and $Y$ be (alpha/quasi)-normed linear spaces over $\mathbb{K}$ and let $T\colon X\rightarrow Y$ be a bounded linear mapping defined on $X$ and taking values in $Y$. Then:
  1. 1.
    $T$ is compact if and only if for every bounded $A\subset X$ the set $T(A)$ is relatively compact in $Y$.
  2. 2.
    If $T$ is compact, then $\alpha(T)=0$.
  3. 3.
    If $Y$ is a complete space and $\alpha(T)=0$, then $T$ is compact.
Proof.
We will gradually prove each statement from the proposition.
  1. 1.
    $\Leftarrow$”: Let $A\subset X$ be bounded and let $\{y_{n}\}$ be a sequence in $T(A)$. Then for every $n\in\mathbb{N}$ there is $x_{n}\in A$ such that $y_{n}=T(x_{n})$. Then $\{x_{n}\}$ is a bounded sequence in $A$ and $\{y_{n}\}=\{T(y_{n})\}$ has a convergent subsequence owing to the compactness of $T$. It follows that $T(A)$ is relatively compact.
    $\Rightarrow$”: Let $A\subset X$ be bounded and and let $\{x_{n}\}$ be a sequence in $A$. Because $\{T(x_{n})\}\subset T(A)$, there exists an increasing sequence of indices $\{n_{k}\}$ such that $\{T(x_{n_{k}})\}$ converges thanks to the relative compactness of $T(A)$. So $T$ satisfies the definition of a compact operator.
  2. 2.
    Let T be compact. From (i) we have that for every bounded $A\subset X$ the set $T(A)$ is relatively compact in $Y$ therefore $T(B_{X})$ is relatively compact. Hence, $T(B_{X})$ is totally bounded. Consequently, for every $\varepsilon>0$, one has $T(B_{X})\subset\bigcup_{i=1}^{m}(y_{i}+\varepsilon B_{Y})$, in which $y_{i}\in Y$ for $i\in\{1,\dots,m\}$, and $m\in\mathbb{N}$. Therefore, $\alpha(T)=0$.
  3. 3.
    Since $\alpha(T)=0$, for every $\varepsilon>0$ we have $T(B_{X})\subset\bigcup_{i=1}^{m}(y_{i}+\varepsilon B_{Y})$, where $y_{i}\in Y$ and $m\in\mathbb{N}$, which implies that $T(B_{X})$ is totally bounded. Because $Y$ is complete, this implies that $T(B_{X})$ is relatively compact in $Y$.
    Now let $A\subset X$ be a bounded set and let $\{x_{n}\}\subset A$ be a sequence from $A$. Then there exists $r>0$ such that $\{x_{n}\}\subset A\subset B_{X}(0,r)$. Then $\frac{1}{r}x_{n}\subset B_{X}$. Now, similarly to the proof of (i), $\{T(\frac{1}{r}x_{n})\}\subset T(B_{X})$, and there exists an increasing sequence of indices $\{n_{k}\}$ such that $\{T(\frac{1}{r}x_{n_{k}})\}$ converges owing to the relative compactness of $T(B_{X})$. But then also, by the linearity of $T$ and linearity of limits, we get that $\{T(x_{n})\}=\{rT(\frac{1}{r}x_{n_{k}})\}$ converges. Therefore, for every bounded $A\subset X$ the set $T(A)$ is relatively compact in $Y$. So, from 1., $T$ is compact.
Definition 1.5 (Embedding).
Assume that $(\ell,\|\cdot\|_{\ell})$ and $(w,\|\cdot\|_{w})$ are sequence (alpha/quasi)-normed linear spaces over $\mathbb{K}$. We say that $\ell$ is a embedded into $w$ (we shall denote this as $\ell\hookrightarrow w$) if $\ell$ is a linear subspace of $w$ and the identity operator is bounded in the sense that there exists a constant $C\geq 0$ such that for every $a\in\ell$ one has $a\in w$ and $\|a\|_{w}\leq C\|a\|_{\ell}.$
We denote by $\mathcal{P}(\mathbb{N})$ the power set of the set $\mathbb{N}$ of all natural numbers. We endow the set $\mathbb{N}$ with the usual $\sigma$-finite counting measure $m\colon\mathcal{P}(\mathbb{N})\to[0,\infty]$.
Below we define a decreasing rearrangement for sequences. We use the definition from the book [2, Definition 1.1 and Definition 1.5; page 36–39] and reformulate it only for the special case where we set the space $(X,\mathcal{A},\mu)$ with a $\sigma$-finite measure $\mu$ equal to $(\mathbb{N},\mathcal{P}(\mathbb{N}),m)$.
Definition 1.6 (Distribution function and decreasing rearrangement).
Let us consider the measure space $(\mathbb{N},\mathcal{P}(\mathbb{N}),m)$. For a sequence $\{a\}_{n=1}^{\infty}\subset\mathbb{K}$ we define its distribution function $m_{a}\colon[0,\infty)\rightarrow[0,\infty]$ in the following way:
$m_{a}(\omega)=m\{n\in\mathbb{N}\colon|a_{n}|>\omega\}.$
Then for $t\geq 0$ we define the decreasing rearrangement:
$a^{*}(t)=\inf\{\omega>0\colon m_{a}(\omega)\leq t\}$
and the decreasing rearrangement $\{a^{*}\}_{n=1}^{\infty}$ of the sequence $\{a\}_{n=1}^{\infty}$ as
$a_{n}^{*}=a^{*}(t)=\inf\{\omega>0\colon m_{a}(\omega)\leq n-1\}$
for $n-1\leq t<n$.
When equipped with the decreasing rearrangement, we are in a position to introduce the crucial notion of a rearrangement-invariant lattice, which will have decisive significance for determining maximal noncompactness of an embedding into the space $c_{0}$.
Definition 1.7 (Rearrangement-invariant lattice).
Let $(\ell,\|\cdot\|_{\ell})$ be a
(alpha/quasi)-normed sequence space over $\mathbb{K}$. We say that $\ell$ is a rearrangement-invariant lattice if it satisfies the following axioms:
  1. 1.
    Let $b\in\ell$ and let $a$ be a sequence such that for all $n\in\mathbb{N}\colon|a_{n}|\leq|b_{n}|$, then also $a\in\ell$ and $\|a\|_{\ell}\leq\|b\|_{\ell}$.
  2. 2.
[…]
We now introduce a notion which, as we will show later, will be important in determining whether an embedding operator is not maximally non-compact.
Definition 1.8 (Sequence space span).
Let $\ell$ be a (alpha/quasi)-normed sequence space over $\mathbb{K}$. We define the span $\sigma_{\ell}$ of $\ell$ as follows:
(1.1) $\sigma_{\ell}:=\sup_{y\in B_{\ell}}\left(\sup_{n\in\mathbb{N}}y_{n}-\inf_{n\in\mathbb{N}}y_{n}\right).$
We will recall the definitions of some familiar sequence spaces, namely, $\ell_{p}$ for $1\leq p<\infty$ and $c_{0}$. One has
$\ell_{p}=\{\{x_{n}\}_{n=1}^{\infty}\subset\mathbb{K};\sum_{n=1}^{\infty}|x_{n}|^{p}<+\infty\}\text{ with norm }\|x\|_{p}=\left(\sum_{n=1}^{\infty}|x_{n}|^{p}\right)^{1/p},$
$\ell_{\infty}=\{\{x_{n}\}_{n=1}^{\infty}\subset\mathbb{K};\sup_{n\in\mathbb{N}}|x_{n}|<+\infty\}\text{ with norm }\|x\|_{\infty}=\sup_{n\in\mathbb{N}}|x_{n}|,$
$c_{0}=\{\{x_{n}\}_{n=1}^{\infty}\subset\mathbb{K};\;\lim_{n\to\infty}x_{n}=0\}\text{ with norm }\|x\|_{\infty}.$
We shall now introduce the Lorentz sequence spaces, a pivotal example of a scale of sequence spaces whose rich and diverse structure will enable us to illustrate on them our general results.
Definition 1.9 (Lorentz sequence spaces $\ell^{p,q}$).
Let $p,q\in(0,\infty]$. We define Lorentz sequence space $\ell^{p,q}$ as a space of all sequences $a=\{a_{n}\}_{n=1}^{\infty}\subset\mathbb{K}$, for which the value $\|a\|_{p,q}$ defined below is finite:
$\|a\|_{p,q}=\begin{cases}\left(\displaystyle\sum_{n=1}^{\infty}(a_{n}^{*})^{q}n^{\frac{q}{p}-1}\right)^{\frac{1}{q}},&\text{if $0<q<\infty$ and $0<p\leq\infty$},\\ \\ \displaystyle\sup_{n\in\mathbb{N}}\{n^{1/p}a_{n}^{*}\},&\text{if $q=\infty$ and $0<p\leq\infty$},\end{cases}$
where $a^{*}=\{a_{n}^{*}\}_{n=1}^{\infty}$ is the decreasing rearrangement of $a$.
Remark 1.10.
Throughout, we adopt the convention $1/\infty=0$.
Remark 1.11.
Below are some of the basic properties of Lorentz sequence spaces which follow directly from the definition:
  1. (1)
    $\|a\|_{p,q}=\left(\sum_{n=1}^{\infty}(a_{n}^{*})^{q}n^{\frac{q}{p}-1}\right)^{\frac{1}{q}}=\|n^{\frac{1}{p}-\frac{1}{q}}a^{*}\|_{\ell^{q}}$,
  2. (2)
    $\ell^{p,p}=\ell^{p}$.
Remark 1.12.
If $1\leq q\leq p\leq\infty$, then $\|\cdot\|_{p,q}$ is a norm. If $0<p<q\leq\infty$, $\|\cdot\|_{p,q}$ is not a norm but it is a quasi-norm satisfying, for any $a,b\in\ell^{p,q}$,
$\|a+b\|_{p,q}\leq 2^{1/p}(\|a\|_{p,q}+\|b\|_{p,q}).$
The proof can be found in [10, p. 76, Proposition 1].
Remark 1.13.
Any Lorentz sequence space $\ell^{p,q}$, $p,q\in(0,\infty]$, trivially satisfies the axioms in Definition 1.7. Therefore, it is a rearrangement-invariant lattice.
We will now formulate two elementary observations which will come handy in the sequel. The statements are known but, for the reader’s convenience, we recall their simple verifications.
Lemma 1.14.
Let $a,b>0$ and $q\geq 1$, then
$(a+b)^{q}\leq 2^{q-1}(a^{q}+b^{q}).$
Proof.
We know that the function $f(t)=t^{q}$ is convex for $q\geq 1$, so $f(\lambda t+(1-\lambda)t)\leq\lambda f(t)+(1-\lambda)f(t)$. From this it already follows that
$\left(\frac{1}{2}\right)^{q}(a+b)^{q}=\left(\frac{1}{2}a+\frac{1}{2}b\right)^{q}\leq\frac{1}{2}(a^{q}+b^{q})$
and we are done. ∎
Lemma 1.15.
Let $a,b>0$ and $0<q<1$, then
$(a+b)^{q}\leq a^{q}+b^{q}.$
Proof.
It is equivalent to show
$\left(\frac{a}{b}\right)^{q}+1\geq\left(\frac{a}{b}+1\right)^{q}$
We will prove this by showing that the function
$f(x)=x^{q}+1-\left(x+1\right)^{q}$
is nonnegative and increasing for $x\geq 0$. We have $f(0)=0$ and
$f^{\prime}(x)=qx^{q-1}-q\left(x+1\right)^{q-1}.$
Then $f^{\prime}(x)\geq 0$ for all $x\geq 0$, because q1<0q-1<0 and $x^{q-1}\geq\left(x+1\right)^{q-1}$. ∎
2. Main results
We begin with a sufficient condition for an embedding between fairly general sequence spaces to be maximally noncompact. For the key idea of our approach to work it is important that the spaces are contained by $c_{0}$.
Proposition 2.1.
Let $\ell$ and $w$ be sequence (alpha/quasi)-normed linear spaces over $\mathbb{K}$ satisfying:
  1. (1)
    $\ell,\ w\subset c_{0}$,
  2. (2)
    for all $x\in w$ and for all $k\in\mathbb{N}$ it holds that $|x_{k}|\leq\|x\|_{w}$,
  3. (3)
    the embedding $I\colon\ell\rightarrow w$ satisfies $0<\|I\|\leq 1$, and $e^{j}\in B_{\ell}$ for every $j\in\mathbb{N}$, where $\|I\|=\displaystyle\sup_{x\in B_{\ell}}\|x\|_{w}$.
Then $I$ is maximally non-compact.
Proof.
We will prove the proposition by contradiction. Let us assume that $I$ is not maximally non-compact. Then $\alpha(I)<\|I\|$ and we take $\varrho\in(\alpha(I),\|I\|)$. Then from the definition of the ball measure of non-compactness we find $m\in\mathbb{N}$ and $y^{1},\dots,y^{m}\in w$ such that:
$I(B_{\ell})\subset\bigcup_{i=1}^{m}(y^{i}+\varrho B_{w}).$
Note that $y^{i}\in w\subset c_{0}$ for each $i\in\{1,\dots,m\}$. So owing to the definition of a limit we can find $j_{0}\in\mathbb{N}$ satisfying: $|(y^{i})_{j}|<\|I\|-\varrho$ for all $j\geq j_{0}$ and for all $i\in\{1,\dots,m\}$. Now, for every $j\geq j_{0}$, the sequence $\|I\|e^{j}=(0,\dots,\ 0,\ \|I\|,\ 0,\dots)$ satisfies $I||e^{j}\in I(B_{\ell})$, but $\|I\|e^{j}\notin(y^{i}+\varrho B_{w})$ for any $i\in\{1,\dots,m\}$. Because for every $z\in(y^{i}+\varrho B_{w})$, we have $(z)_{j}<(\|I\|-\varrho)+\varrho=\|I\|$, so we get our contradiction, hence $\alpha(I)=\|I\|$, and $I$ is maximally non-compact. ∎
We will now present an example to show that the Proposition 2.1 does not hold without the second condition.
Example 2.2.
Let $p\in[1,\infty)$, let $\ell=\ell_{p}$ and let $w$ be a weighted $\ell_{p}$ space endowed with the norm
$\|x\|_{w}=\left(|x_{1}|^{p}+\frac{1}{2}\sum_{n=2}^{\infty}|x_{n}|^{p}\right)^{1/p}.$
Then $\ell$, $w$ and embedding $I\colon\ell\rightarrow w$ satisfy all but the second assumption of Proposition 2.1, but $I$ is not maximally non-compact and $\alpha(I)\leq\left(\frac{1}{2}\right)^{\frac{1}{p}}<\|I\|$.
Indeed, this can be verified by the following reasoning.
  1. 1.
    One has $\ell,w\subset c_{0}$ from the necessary condition for the sum convergence.
  2. 2.
    The space $w$ does not satisfy the condition that for all $x\in w$ and for all $k\in\mathbb{N}$ it holds that $|x_{k}|\leq\|x\|_{w}$. For example take $x=e^{2}=(0,1,0,0,\dots)$, then $1=|x_{2}|>\|x\|_{w}=\left(\frac{1}{2}\right)^{\frac{1}{p}}$.
  3. 3.
    We will verify the third assumption of the Proposition 2.1.
    1. a)
      It is easy to see that $e^{j}\in B_{\ell}$ for every $j\in\mathbb{N}$.
    2. b)
      For all $x\in\ell_{p}$ we have $\|x\|_{w}\leq\|x\|_{p}$.
    Thus, $\|I\|\leq 1$ for any $p\in[1,\infty)$. Now we will show that $I$ achieves $1$ on the closed unit ball. For $e^{1}\in B_{\ell_{p}}$ we have $\|e^{1}\|_{p}=\|e^{1}\|_{w}=1$, hence $\|I\|=1$.
We show that $\alpha(I)\leq\left(\frac{1}{2}\right)^{\frac{1}{p}}$. First we notice that for $n>1$ each member $y_{n}$ of the sequence $y\in B_{w}$ can be $2^{\frac{1}{p}}$ times larger than in $B_{\ell_{p}}$. Therefore, the ball $\left(\frac{1}{2}\right)^{\frac{1}{p}}B_{w}$ would cover all sequences $x\in B_{\ell_{p}}$ such that $x_{1}=0$. The remaining sequences $x\in B_{\ell_{p}}$, where $0<x_{1}\leq 1$ can be covered with balls $\pm e^{1}+\left(\frac{1}{2}\right)^{\frac{1}{p}}B_{w}$ and $\pm\left(\frac{1}{2}\right)^{\frac{1}{p}}e^{1}+\left(\frac{1}{2}\right)^{\frac{1}{p}}B_{w}$. Altogether, we showed that the radius in the definition of maximal non-compactness is less than or equal to $\left(\frac{1}{2}\right)^{\frac{1}{p}}$. Hence $I$ is not maximally non/compact.
We shall now state our first main result, in which we formulate a sufficient condition for the maximal non-compactness of an embedding of rearrangement-invariant lattices contained in $c_{0}$.
Theorem 2.3.
Let $\ell$ and $w$ be sequence (alpha/quasi)-normed linear spaces over $\mathbb{K}$ satisfying:
  1. (1)
    $\ell\hookrightarrow w\subset c_{0}$,
  2. (2)
    $\ell,\ w$ are rearrangement-invariant lattices,
  3. (3)
    the embedding operator $I\colon\ell\rightarrow w$ satisfies $0<\|I\|<\infty$, where $\|I\|=\displaystyle\sup_{x\in B_{\ell}}\|x\|_{w}$.
Then the embedding operator $I$ is maximally non-compact.
Proof.
We will prove the proposition by contradiction. Let us assume that $I$ is not maximally non-compact, then $\alpha(I)<\|I\|$. We take $\varrho\in(\alpha(I),\|I\|)$ and find $\lambda\in(0,1)$ such that:
(2.1) $\frac{\varrho}{1-\lambda}<\|I\|.$
From the definition of $\|I\|$ we find $x\in B_{\ell}$ satisfying:
(2.2) $\|x\|_{w}>\frac{\varrho}{1-\lambda}.$
Then from the definition of the ball measure of non-compactness we find $m\in\mathbb{N}$ and $y^{1},\dots,y^{m}\in w$ such that
(2.3) $I(B_{\ell})\subset\bigcup_{i=1}^{m}(y^{i}+\varrho B_{w}).$
Let us now define a new sequence $\varepsilon=\{\varepsilon_{k}\}_{k=1}^{\infty}$. For all $k\in\mathbb{N}$ we set
(2.4) $\varepsilon_{k}=\lambda x_{k}^{*}.$
Owing to (2.4), $\varepsilon\in c_{0}$. Then from the positive homogeneity of (alpha/quasi)-norm, our prerequisite that $w$ is a rearrangement-invariant lattice, and inequality (2.2), we get
(2.5) $\|x^{*}-\varepsilon\|_{w}=\|\{x_{k}^{*}-\lambda x_{k}^{*}\}\|_{w}=\|(1-\lambda)x^{*}\|_{w}=(1-\lambda)\|x\|_{w}>\varrho.$
Now $y^{i}\in w\subset c_{0}$ for each $i\in\{1,\dots,m\}$. So for every $k\in\mathbb{N}$, there exists $j_{k}\in\mathbb{N}$, such that for every $k\geq 2$, one has $j_{k+1}>j_{k}$, and for all $i\in\{1,\dots,m\}$
(2.6) $|(y^{i})_{j_{k}}|\leq\varepsilon_{k}.$
Let us now define a new sequence:
(2.7) $a:=\sum_{k=1}^{\infty}x_{k}^{*}e^{j_{k}}.$
Then $a^{*}=x^{*}$. Consequently, $a\in B_{\ell}$, $\|a\|_{\ell}=\|x\|_{\ell}$ and $\|a\|_{w}=\|x\|_{w}$, since $\ell$ and $w$ are rearrangement-invariant lattices and $x\in B_{\ell}$. Now for every $i\in\{1,\dots,m\}$ and every $k\in\mathbb{N}$ from inequality (2.6) and definition of $a$:
(2.8) $a_{j_{k}}-y_{j_{k}}^{i}=x_{k}^{*}-y_{j_{k}}^{i}\geq x_{k}^{*}-\varepsilon_{k}\geq 0.$
This together with inequality (2.5) yields
$\|a-y^{i}\|_{w}$ $=\displaystyle\|\sum_{k=1}^{\infty}e^{j_{k}}x^{*}_{k}-y^{i}\|_{w}\geq\displaystyle\|\sum_{k=1}^{\infty}e^{j_{k}}(x^{*}_{k}-y_{j_{k}}^{i})\|_{w}$
$\geq\displaystyle\|\sum_{k=1}^{\infty}e^{j_{k}}(x_{k}^{*}-\varepsilon_{k})\|_{w}$
$=\|x^{*}-\varepsilon\|_{w}$
$>\varrho.$
We showed that $a\notin(y^{i}+\varrho B_{w})$ for all $i\in\{1,\dots,m\}$. Which gives us a contradiction with our claim (2.3) for any choice of $\varrho\in(\alpha(I),\|I\|)$. Therefore $I$ is maximally non-compact, and $\alpha(I)=\|I\|$. ∎
Our next main result gives a general sufficient condition for an embedding operator into $\ell^{\infty}$ to avoid maximal non-compactness.
Theorem 2.4.
Let $\ell$ be a sequence (alpha/quasi)-normed linear space over $\mathbb{K}$ satisfying $\ell\hookrightarrow\ell^{\infty}$ and
(2.9) $\|I\|\leq\sigma_{\ell}<2\|I\|.$
Then the embedding operator $I\colon\ell\rightarrow\ell^{\infty}$ is not maximally non-compact. Moreover, $\alpha(I)\leq\sigma_{\ell}/2$.
Proof.
Denote $\varrho\in(\sigma_{\ell}/2,\|I\|)$ and consider $m\in\mathbb{N}$ such that
(2.10) $\left(1+\frac{1}{m}\right)\frac{\sigma_{\ell}}{2}<\varrho.$
Define $\lambda_{k}=\frac{\sigma_{\ell}k}{2m}$ for $k=-m,\dots,m$ and let $y^{k}$ be a constant sequence defined by $(y^{k})_{j}=\lambda_{k}$ for every $j\in\mathbb{N}$ and $k=-m,\dots,m$. We will show that
(2.11) $B_{\ell}\subset\bigcup_{i=-m}^{m}(y^{i}+\varrho B_{\ell^{\infty}}),$
proving $\alpha(I)\leq\varrho<\|I\|$. Assume $y\in B_{\ell}$. Then $y\in B_{\ell^{\infty}}(0,\|I\|)$ and $|y_{j}|\leq\|I\|\leq\sigma_{\ell}$ for every $j\in\mathbb{N}$. Now from (1.1) it follows that $\sup y-\inf y\leq\sigma_{\ell}$. We shall now distinguish three cases.
  1. a)
    If $\inf y=-\sigma_{\ell}$, then $\sup y\in[-\sigma_{\ell},0]$. Thus, $y_{j}\in[-\sigma_{\ell},0]$ for each $j\in\mathbb{N}$, and we claim that $y\in y^{-m}+\varrho B_{\ell^{\infty}}$. Indeed, since $\varrho>\frac{\sigma_{\ell}}{2}$ and for every $j\in\mathbb{N}$ we have $(y^{-m})_{j}=\lambda_{-m}=\frac{\sigma_{\ell}(-m)}{2m}=-\frac{\sigma_{\ell}}{2}$, the claim follows.
  2. b)
    If $\inf y\in(-\sigma_{\ell},0]$, then there is a unique $k\in\{-m+1,\dots,m\}$ such that $\inf y+\sigma_{\ell}/2\in(\lambda_{k-1},\lambda_{k}]=(\frac{\sigma_{\ell}(k-1)}{2m},\frac{\sigma_{\ell}k}{2m}]\subset(-\sigma_{\ell}/2,\sigma_{\ell}/2]$. Then by the choice of $\varrho>\sigma_{\ell}/2$ and inequality (2.9),
    $\lambda_{k}+\varrho>\lambda_{k}+\sigma_{\ell}/2\geq\inf y+\sigma_{\ell}\geq\sup y.$
    Here, we get the second inequality from infy + σ_ℓ ∈(λ_k-1 + σ_ℓ/2, λ_k + σ_ℓ/2]. On the other hand, using the definition of $\lambda_{k}$ and inequality (2.10), we arrive at
    $\inf y>\lambda_{k-1}-\sigma_{\ell}/2=\frac{\sigma_{\ell}k-\sigma_{\ell}-\sigma_{\ell}m}{2m}=\lambda_{k}-\left(1+\frac{1}{m}\right)\frac{\sigma_{\ell}}{2}>\lambda_{k}-\varrho.$
  3. c)
    If $\inf y\in(0,\sigma_{\ell}]$, then $y_{j}\in[0,\sigma_{\ell}]$ for each $j\in\mathbb{N}$, hence, evidently, $y\in y^{m}+\varrho B_{\ell^{\infty}}$.
Altogether, we showed that $y\in y^{k}+\varrho B_{\ell^{\infty}}$, and (2.11) follows. Therefore $\alpha(I)\leq\sigma_{\ell}/2<\|I\|$ and $I$ is not maximally non-compact. ∎
We will now state an example of a maximally non-compact embedding, where the target space is not a subspace of $c_{0}$, and therefore Theorem 2.3 cannot be used.
Theorem 2.5.
Let $I$ be the embedding operator $I\colon c_{0}\rightarrow\ell_{\infty}$. Then $I$ is maximally non-compact.
Proof.
First, we compute the operator norm of $I$. Let $x=\{x_{n}\}_{n=1}^{\infty}\in c_{0}$, then $\|I(x)\|=\|x\|_{\infty}$. So $I$ is a linear isometry from $c_{0}$ to $\ell_{\infty}$, which implies that $\|I\|=1$ and it achieves its norm on an arbitrary element $x\in B_{c_{0}}$ such that $\|x\|_{\infty}=1$.
We will proceed with the proof by contradiction. Let us assume that $I$ is not maximally non-compact. Then $\alpha(I)<\|I\|=1$ and we take $\varrho\in(\alpha(I),1)$. Then from the definition of the ball measure of non-compactness we find $y^{1},\dots,y^{m}\in\ell_{\infty}$, where $m\in\mathbb{N}$, such that:
$B_{c_{0}}\subset\bigcup_{i=1}^{m}(y^{i}+\varrho B_{\ell_{\infty}}).$
We will define a sequence $a=(a_{j})_{j=1}^{\infty}$ in the following way:
$a_{j}=\begin{cases}1,&\text{if $(y^{j})_{j}<0$ and $j\in\{1,\dots,m\}$},\\ -1,&\text{if $(y^{j})_{j}\geq 0$ and $j\in\{1,\dots,m\}$},\\ 0,&\text{for $j>m$}.\end{cases}$
Trivially $a\in c_{0}$, because it has only a finite number of non-zero elements. Also $\|a\|_{\infty}=\sup_{n\in\mathbb{N}}|a_{n}|=1$. So $a\in B_{c_{0}}$.
We will show that $a\notin y^{i}+\varrho B_{\ell_{\infty}}$ for any $i\in\{1,\dots,m\}$. Fix $i\in\{1,\dots,m\}$ and $j\in\mathbb{N}$. Then for all $z\in y^{i}+\varrho B_{\ell_{\infty}}$ it holds that $z_{j}\in[(y^{i})_{j}-\varrho,\ (y^{i})_{j}+\varrho]$. We have $\varrho\in(0,1)$, thus for any $j\in\mathbb{N}$ it cannot happen that $\{1,-1\}\subset[(y^{i})_{j}-\varrho,\ (y^{i})_{j}+\varrho]$, because this interval can contain at most one of the elements of the set $\{1,\ -1\}$. Let us take $j\in\{1,\dots,m\}$ arbitrarily. From the definition of $a$ it follows that the element $a_{j}\notin[(y^{j})_{j}-\varrho,\ (y^{j})_{j}+\varrho]$. So, the whole sequence $a$ cannot belong to the ball $(y^{j}+\varrho B_{\ell_{\infty}})$. This holds for every $j\in\{1,\dots,m\}$. And thus we have shown that $a\notin\bigcup_{i=1}^{m}(y^{i}+\varrho B_{\ell_{\infty}})$ and we get a contradiction. ∎
Remark 2.6.
We could not use Theorem 2.4 for the embedding $I\colon c_{0}\rightarrow\ell_{\infty}$ in Theorem 2.5, because $\sigma_{c_{0}}=2$. It is achieved for example on the sequence $y=(1,-1,0,0,\dots)\in B_{c_{0}}$.
3. Lorentz sequence spaces
In this section we shall focus on one particular class of sequence spaces, namely the two-parameter scale of Lorentz spaces $\ell^{p,q}$. This class of spaces is surprisingly rich and contains many diverse examples of interesting relations.
To begin, we single out those Lorentz spaces, which are contained in $c_{0}$.
Lemma 3.1.
Let $p,q\in(0,\infty]$ and $\min\{p,q\}<\infty$, then $\ell^{p,q}\subset c_{0}$.
Proof.
Let us assume that $\ell^{p,q}\not\subset c_{0}$ and take $a\in\ell^{p,q}\setminus c_{0}$. Then there exists $\varepsilon>0$, such that for all $n_{0}\in\mathbb{N}$ there exists $n\geq n_{0},\ n\in\mathbb{N}\colon|a_{n}|\geq\varepsilon$, which implies that $a_{n}^{*}\ \geq\ \varepsilon$ for $\forall n\in\mathbb{N}$. First, let us look at the case when $q<\infty$ and $p\in(0,\infty]$. Then:
$\infty>\|a\|_{p,q}\geq\varepsilon\left(\sum_{n=1}^{\infty}(n^{\frac{1}{p}-\frac{1}{q}})^{q}\right)^{\frac{1}{q}}=\varepsilon\left(\sum_{n=1}^{\infty}n^{\frac{q}{p}-1}\right)^{\frac{1}{q}}=\infty,$
because for all $p\in(0,\infty]$ one has $(\frac{q}{p}-1)\geq-1$, and we know that $\sum_{n=1}^{\infty}n^{\alpha}$ diverges for $\alpha\geq-1$. This gives us a contradiction with our assumption that $a\in\ell^{p,q}$. Similarly, for $q=\infty$ and $p\in(0,\infty)$ we get $\infty>\|a\|_{p,\infty}\geq\sup_{n\in\mathbb{N}}\{\varepsilon n^{1/p}\}=\infty$, which is a contradiction once again. ∎
The statements of the following two lemmas occur in the article [10, p. 77-78]. Since, however, the verification of the statements is only sketched there, we shall include detailed proofs.
Lemma 3.2.
Let $p\in(0,\infty]$, $q\in(0,\infty)$ and $a=\{a_{n}\}_{n=1}^{\infty}\in\ell_{p,q}$, then for all $n\in\mathbb{N}$ it holds that
  1. 1.
    $a_{n}^{*}\leq\left(\frac{q}{p}\right)^{\frac{1}{q}}n^{-\frac{1}{p}}\|a\|_{p,q}$, if $0<p\leq q<\infty$,
  2. 2.
    $a_{n}^{*}\leq n^{-\frac{1}{p}}\|a\|_{p,q}$, if $0<q<p\leq\infty$.
Proof.
We will prove the individual cases:
  1. 1.
    Let $0<p\leq q<\infty$ and fix $n\in\mathbb{N}$, Then:
    $\|a\|_{p,q}^{q}$ $=\displaystyle\sum_{i=1}^{\infty}i^{\frac{q}{p}-1}(a_{i}^{*})^{q}$
    $\geq\displaystyle\sum_{i=1}^{n}i^{\frac{q}{p}-1}(a_{i}^{*})^{q}$
    $\geq(a_{n}^{*})^{q}\displaystyle\sum_{i=1}^{n}i^{\frac{q}{p}-1}$
    $\geq(a_{n}^{*})^{q}\frac{p}{q}\displaystyle\sum_{i=1}^{n}\left(i^{\frac{q}{p}}-(i-1)^{\frac{q}{p}}\right)$
    $=(a_{n}^{*})^{q}\left(\frac{p}{q}\right)n^{\frac{q}{p}}.$
    Here, the second inequality follows from the fact that $(a_{i}^{*})$ is the decreasing rearrangement. The third inequality results from the Lagrange mean value theorem. Indeed, for all $i\in\mathbb{N}$, there exists $\xi\in(i-1,i)$, that satisfies for the function $f(t)=\frac{p}{q}t^{\frac{q}{p}}$, where $t\geq 0$ following equality: f^’(ξ) = ξ^qp - 1 = pq[i^qp-(i-1)^qp]. Also $i^{\frac{q}{p}-1}\geq\xi^{\frac{q}{p}-1}$, because $\frac{q}{p}\geq 1$ and $i>\xi>1$. From the estimates above we get the third inequality and 1. is proven.
  2. 2.
    The assertion can be shown analogously to the one in 1. We get the following inequality:
    $\|a\|_{p,q}^{q}$ $\geq(a_{n}^{*})^{q}\displaystyle\sum_{i=1}^{n}i^{\frac{q}{p}-1}$
    $\geq(a_{n}^{*})^{q}nn^{\frac{q}{p}-1}$
    $=(a_{n}^{*})^{q}n^{\frac{q}{p}}.$
    Here the second inequality follows from the fact that for all $i\leq n$ it holds that $\frac{q}{p}\in[0,1)$. Therefore $\left(\frac{q}{p}-1\right)\in[-1,0)$, so we get $n^{\left(\frac{q}{p}-1\right)}\leq i^{\left(\frac{q}{p}-1\right)}$, and 2. is proven.
Theorem 3.3.
Let $0<p\leq\infty$, $0<q_{1}<q_{2}\leq\infty$, Then:
$\ell_{p,q_{1}}\hookrightarrow\ell_{p,q_{2}},$
and for all $a\in\ell_{p,q_{1}}$ it holds that
  1. 1.
    $\|a\|_{p,q_{2}}\leq\left(\frac{q_{1}}{p}\right)^{\frac{1}{q_{1}}-\frac{1}{q_{2}}}\|a\|_{p,q_{1}}$; if $p<q_{1}$,
  2. 2.
    $\|a\|_{p,q_{2}}\leq\|a\|_{p,q_{1}}$; if $p\geq q_{1}$.
Proof.
We will treat each case separately.
  1. 1.
    1. (a)
      Let $q_{2}<\infty$. From the first part of Lemma 3.2 we get:
      $\|a\|_{p,q_{2}}^{q_{2}}$ $=\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p}-1}(a_{n}^{*})^{q_{2}-q_{1}}(a_{n}^{*})^{q_{1}}$
      $\leq\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p}-1}\left(\frac{q_{1}}{p}\right)^{\frac{q_{2}-q_{1}}{q_{1}}}n^{-\frac{q_{2}-q_{1}}{p}}\|a\|_{p,q_{1}}^{q_{2}-q_{1}}(a_{n}^{*})^{q_{1}}$
      $=\left(\frac{q_{1}}{p}\right)^{\frac{q_{2}}{q_{1}}-1}\|a\|_{p,q_{1}}^{q_{2}-q_{1}}\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{1}}{p}-1}(a_{n}^{*})^{q_{1}}$
      $=\left(\frac{q_{1}}{p}\right)^{\frac{q_{2}}{q_{1}}-1}\|a\|_{p,q_{1}}^{q_{2}-q_{1}}\|a\|_{p,q_{1}}^{q_{1}}$
      $=\left(\frac{q_{1}}{p}\right)^{\frac{q_{2}}{q_{1}}-1}\|a\|_{p,q_{1}}^{q_{2}}.$
    2. (b)
      Now, let $q_{2}=\infty$:
      $\|a\|_{p,q_{2}}$ $=\displaystyle\sup_{n\in\mathbb{N}}n^{\frac{1}{p}}a_{n}^{*}$
      $\leq\displaystyle\sup_{n\in\mathbb{N}}n^{\frac{1}{p}}\left(\frac{q_{1}}{p}\right)^{\frac{1}{q_{1}}}n^{-\frac{1}{p}}\|a\|_{p,q_{1}}$
      $=\left(\frac{q_{1}}{p}\right)^{\frac{1}{q_{1}}}\|a\|_{p,q_{1}}.$
      Here the first inequality again follows from Lemma 3.2.
    So 1. is proven.
  2. 2.
    Now let $p\geq q_{1}$. Then, the following inequality follows from the second part of Lemma 3.2:
    $\|a\|_{p,q_{2}}^{q_{2}}$ $=\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p}-1}(a_{n}^{*})^{q_{2}-q_{1}}(a_{n}^{*})^{q_{1}}$
    $\leq\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p}-1}n^{-\frac{q_{2}-q_{1}}{p}}\|a\|_{p,q_{1}}^{q_{2}-q_{1}}(a_{n}^{*})^{q_{1}}$
    $=\|a\|_{p,q_{1}}^{q_{2}-q_{1}}\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{1}}{p}-1}(a_{n}^{*})^{q_{1}}$
    $=\|a\|_{p,q_{1}}^{q_{2}-q_{1}}\|a\|_{p,q_{1}}^{q_{1}}$
    $=\|a\|_{p,q_{1}}^{q_{2}}.$
We will now concentrate on establishing the embeddings $\ell_{p_{1},q_{1}}\hookrightarrow\ell_{p_{2},q_{2}}$ between the Lorentz sequence spaces in dependence on the values of $p_{1},p_{2}$ and $q_{1},q_{2}$. Then we will try to calculate the exact value of the operator norms of these embeddings.
Theorem 3.4.
Let $0<p_{1}<p_{2}\leq\infty$ and $q_{1},q_{2}\in(0,\infty]$. Then
$\ell_{p_{1},q_{1}}\hookrightarrow\ell_{p_{2},q_{2}},$
and for all $a\in\ell_{p_{1},q_{1}}$ holds:
  1. 1.
    $\|a\|_{p_{2},q_{2}}\leq\|a\|_{p_{1},q_{1}}$, if $q_{1}\leq p_{1}$,
  2. 2.
    $\|a\|_{p_{2},q_{2}}\leq\left(\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p_{2}}-\frac{q_{2}}{p_{1}}-1}\right)^{1/q_{2}}\|a\|_{p_{1},\infty}$, if $p_{1}<q_{1}=\infty$ and $q_{2}<\infty$,
  3. 3.
    $\|a\|_{p_{2},q_{2}}\leq\left(\frac{q_{1}}{p_{1}}\right)^{\frac{1}{q_{1}}-\frac{1}{q_{2}}}\|a\|_{p_{1},q_{1}}$, if $p_{1}<q_{1}<q_{2}\leq\infty$,
  4. 4.
    $\|a\|_{p_{2},q_{2}}\leq\|a\|_{p_{1},q_{1}}$, if $p_{1}<q_{1}<\infty$ and $q_{1}\geq q_{2}$.
  5. 5.
    $\|a\|_{p_{2},q_{2}}\leq\|a\|_{p_{1},q_{1}}$, if $q_{1}=q_{2}=\infty$.
Proof.
Let $0<p_{1}<p_{2}\leq\infty$. We will prove the individual cases.
  1. 1.
    Let $q_{1}\leq p_{1}<\infty$.
    1. a)
      Let $q_{2}<\infty$. Then
      $\|a\|_{p_{2},q_{2}}^{q_{2}}$ $=\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p_{2}}-1}(a_{n}^{*})^{q_{2}}$
      $=\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p_{2}}-1}(a_{n}^{*})^{q_{2}-q_{1}}(a_{n}^{*})^{q_{1}}$
      $\leq\|a\|_{p_{1},q_{1}}^{q_{2}-q_{1}}\displaystyle\sum_{n=1}^{\infty}n^{-\frac{q_{2}-q_{1}}{p_{1}}}(a_{n}^{*})^{q_{1}}n^{\frac{q_{2}}{p_{2}}-1}$
      $=\|a\|_{p_{1},q_{1}}^{q_{2}-q_{1}}\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p_{2}}-\frac{q_{2}}{p_{1}}}(a_{n}^{*})^{q_{1}}n^{\frac{q_{1}}{p_{1}}-1}$
      $\leq\|a\|_{p_{1},q_{1}}^{q_{2}-q_{1}}\displaystyle\sum_{n=1}^{\infty}(a_{n}^{*})^{q_{1}}n^{\frac{q_{1}}{p_{1}}-1}$
      $\leq\|a\|_{p_{1},q_{1}}^{q_{2}}.$
      Here the first inequality follows from the second part of Lemma 3.2 a_n^* ≤n^-1p1∥ a ∥_p_1,q_1, and we get the second inequality from the fact that n^q2p2 -q2p1¡1.
    2. b)
      Now let $q_{2}=\infty$ then:
      $\|a\|_{p_{2},\infty}$ $=\displaystyle\sup_{n\in\mathbb{N}}a_{n}^{*}n^{\frac{1}{p_{2}}}$
      $\leq\displaystyle\sup_{n\in\mathbb{N}}n^{\frac{1}{p_{2}}}n^{-\frac{1}{p_{1}}}\|a\|_{p_{1},q_{1}}$
      $=\|a\|_{p_{1},q_{1}}\displaystyle\sup_{n\in\mathbb{N}}n^{\frac{1}{p_{2}}}n^{-\frac{1}{p_{1}}}$
      $\leq\|a\|_{p_{1},q_{1}}.$
      The first inequality follows from the second part of Lemma 3.2 a_n^* ≤n^-1p1∥ a ∥_p_1,q_1. And we get the second inequality from the fact that n^1p2 -1p1 ≤1.
  1. 2.
    Let $p_{1}<q_{1}=\infty$ and $q_{2}<\infty$, then:
    $\|a\|_{p_{2},q_{2}}^{q_{2}}$ $=\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p_{2}}-1}(a_{n}^{*})^{q_{2}}$
    $=\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p_{2}}-\frac{q_{2}}{p_{1}}-1}(a_{n}^{*})^{q_{2}}n^{\frac{q_{2}}{p_{1}}}$
    $\leq\left(\displaystyle\sup_{n\in\mathbb{N}}a_{n}^{*}n^{\frac{1}{p_{1}}}\right)^{q_{2}}\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p_{2}}-\frac{q_{2}}{p_{1}}-1}$
    $=\|a\|_{p_{1},\infty}^{q_{2}}\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p_{2}}-\frac{q_{2}}{p_{1}}-1},$
    in which the sum $\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p_{2}}-\frac{q_{2}}{p_{1}}-1}$ converges, because $\frac{q_{2}}{p_{2}}-\frac{q_{2}}{p_{1}}-1<-1$.
  1. 3.
    Let $p_{1}<q_{1}<\infty$ and $q_{2}<\infty$, then:
    $\|a\|_{p_{2},q_{2}}^{q_{2}}$ $=\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p_{2}}-1}(a_{n}^{*})^{q_{2}}$
    $=\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p_{2}}-1}(a_{n}^{*})^{q_{2}-q_{1}}(a_{n}^{*})^{q_{1}}$
    $\leq\|a\|_{p_{1},q_{1}}^{q_{2}-q_{1}}\left(\frac{q_{1}}{p_{1}}\right)^{\frac{q_{2}-q_{1}}{q_{1}}}\displaystyle\sum_{n=1}^{\infty}n^{-\frac{q_{2}-q_{1}}{p_{1}}}(a_{n}^{*})^{q_{1}}n^{\frac{q_{2}}{p_{2}}-1}$
    $=\|a\|_{p_{1},q_{1}}^{q_{2}-q_{1}}\left(\frac{q_{1}}{p_{1}}\right)^{\frac{q_{2}}{q_{1}}-1}\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p_{2}}-\frac{q_{2}}{p_{1}}}(a_{n}^{*})^{q_{1}}n^{\frac{q_{1}}{p_{1}}-1}$
    $\leq\left(\frac{q_{1}}{p_{1}}\right)^{\frac{q_{2}}{q_{1}}-1}\|a\|_{p_{1},q_{1}}^{q_{2}}.$
    Here the first inequality follows from the first part of Lemma 3.2 a_n^* ≤( q1p1)^1q1n^-1p1∥ a ∥_p_1,q_1 and we get the second inequality from the fact that n^q2p2 -q2p1 ≤1.
  1. 4.
    Let $p_{1}<q_{1}<\infty$ and $q_{1}\geq q_{2}$, then from the computations in the 3. part of the proof we have: ∥ a ∥_p_2,q_2≤( q1p1)^1q1 - 1q2∥ a ∥_p_1,q_1,
    where $\frac{q_{1}}{p_{1}}>1$ because $p_{1}<q_{1}<\infty$. But $\frac{1}{q_{1}}-\frac{1}{q_{2}}\leq 0$ because $q_{1}\geq q_{2}$. So $\left(\frac{q_{1}}{p_{1}}\right)^{\frac{1}{q_{1}}-\frac{1}{q_{2}}}\leq 1$ and $\|a\|_{p_{2},q_{2}}\leq\|a\|_{p_{1},q_{1}}$.
  1. 5.
    Let $q_{1}=q_{2}=\infty$, then:
    $\|a\|_{p_{2},\infty}$ $=\displaystyle\sup_{n\in\mathbb{N}}a_{n}^{*}n^{\frac{1}{p_{2}}}$
    $\leq\displaystyle\sup_{n\in\mathbb{N}}n^{\frac{1}{p_{1}}}a_{n}^{*}$
    $=\|a\|_{p_{1},q_{1}},$
    where the inequality follows from $p_{1}<p_{2}$, hence $n^{\frac{1}{p_{1}}}\geq n^{\frac{1}{p_{2}}}$.
Theorem 3.5.
Let $0<p_{1}\leq p_{2}\leq\infty$ and $q_{1}\in[1,\infty]$ and $q_{2}\in[1,\infty]$, then the norm of the embedding operator $I\colon\ell_{p_{1},q_{1}}\rightarrow\ell_{p_{2},q_{2}}$ is equal to:
  1. 1.
    $\|I\|=1$, if $q_{1}\leq p_{1}$,
  2. 2.
    $\|I\|=1$, if $p_{1}<q_{1}<\infty$, $q_{1}\geq q_{2}$ and $p_{1}<p_{2}$,
  3. 3.
    $\|I\|=1$, if $q_{1}=q_{2}=\infty$ and $p_{1}<p_{2}$,
  4. 4.
    $\|I\|=\left(\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p_{2}}-\frac{q_{2}}{p_{1}}-1}\right)^{1/q_{2}}$, if $p_{1}<q_{1}=\infty$, $q_{2}<\infty$ and $p_{1}<p_{2}$.
Proof.
We will prove the individual cases separately:
  1. 1.
    Let $q_{1}\leq p_{1}$. We get the upper estimate of the operator norm for and $p_{1}<p_{2}$ from the first part of Lemma 3.4 and for and $p_{1}=p_{2}$ from the second part of Lemma 3.3. So $\|I\|\leq 1$. Now we will show that $I$ attains $1$ on the closed unit ball. For arbitrary $e^{j}$ we have that $(e^{j})^{*}=e^{1}$ and: ∥e^j∥_p_1,q_1 =(∑n=1 nq1p1- 1((enj)*)q1)^1q1 = (1^q1p1 - 11^q_1)^1q1 = 1. So $e^{j}\in B_{\ell_{p_{1},q_{1}}}$. Now we need to show that also $\|I(e^{j})\|_{p_{2},q_{2}}=\|e^{j}\|_{p_{2},q_{2}}=1$.
    1. a)
      Let $q_{2}<\infty$. Then $\|e^{j}\|_{p_{2},q_{2}}=1$ follows from the computations above.
    2. b)
      Let $q_{2}=\infty$. Then $\|e^{j}\|_{p_{2},q_{2}}=\displaystyle\sup_{n\in\mathbb{N}}((e_{n}^{j})^{*})n^{\frac{1}{p_{2}}}=1$
    It already follows that $\|I\|=1$.
  2. 2.
    Let $p_{1}<q_{1}<\infty$ and $q_{1}\geq q_{2}$. We get the upper estimate for the operator norm from the fourth part of Lemma 3.4. So $\|I\|\leq 1$. We can show that $I$ attains $1$ on the closed unit ball at an arbitrary $e^{j}$. The computations are same as in the proof of 1. So again $\|I\|=1$.
  3. 3.
    Let $q_{1}=q_{2}=\infty$. We get the upper estimate for the operator norm from the fifth part of Lemma 3.4. So $\|I\|\leq 1$. We can show that $I$ attains $1$ on the closed unit ball at an arbitrary $e^{j}$. The computations are similar as in the proof of 1. ∥e^j∥_p_1,q_1 = supn∈N((enj)*)n1p1 = 1 = ∥ejp2,q2
    So again $\|I\|=1$.
  4. 4.
    Let $p_{1}<q_{1}=\infty$ and $q_{2}<\infty$. Once again we get the upper estimate for the operator norm from the second part of Lemma 3.4.
    Therefore $\|I\|\leq\left(\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p_{2}}-\frac{q_{2}}{p_{1}}-1}\right)^{1/q_{2}}$. Now we will again show that $I$ achieves this upper estimate on the closed unit ball. Let us consider the sequence
[…]
    $\|a\|_{p_{2},q_{2}}$ $=\left(\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p_{2}}-1}(a_{n}^{*})^{q_{2}}\right)^{1/q_{2}}$
    $=\left(\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p_{2}}-1}(n^{-\frac{1}{p_{1}}})^{q_{2}}\right)^{1/q_{2}}$
    $=\left(\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p_{2}}-\frac{q_{2}}{p_{1}}-1}\right)^{1/q_{2}}.$
    So indeed $I$ achieves it’s upper estimate on the closed unit ball and 4. is proven.
Theorem 3.6.
Let $0<p<q<\infty$, then the norm of the embedding operator $I\colon\ell_{p,q}\rightarrow\ell_{p,\infty}$ is:
$\|I\|=\left(\frac{q}{p}\right)^{\frac{1}{q}}.$
Proof.
$\|I\|\leq\left(\frac{q}{p}\right)^{\frac{1}{q}}$ from the first part of Lemma 3.3.
We want to show that there exists a sequence $\{a^{n}\}_{n=1}^{\infty}$ of nonzero elements, where $a^{n}\in\ell_{p,q}$ for all $n\in\mathbb{N}$, such that:
(3.1) $\lim_{n\rightarrow\infty}\frac{\|a^{n}\|_{p,\infty}}{\|a^{n}\|_{p,q}}=\left(\frac{q}{p}\right)^{\frac{1}{q_{1}}}.$
This would already imply that $\|I\|=\left(\frac{q}{p}\right)^{\frac{1}{q_{1}}}$, since we would be able to get arbitrarily close to our upper estimate $\left(\frac{q}{p}\right)^{\frac{1}{q}}$ of the operator norm $\|I\|$.
For every $n\in\mathbb{N}$ set $a^{n}=(1,1,1,\dots,1,0,0\dots)$, where $a_{i}^{n}=1$ for $i\in\{1,2,\dots,n\}$ and $a_{i}^{n}=0$ for $i>n$. Then for every $n\in\mathbb{N}$:
$\|a^{n}\|_{p,\infty}=\sup_{i\leq n}i^{\frac{1}{p}}=n^{\frac{1}{p}},$
while
$\|a^{n}\|_{p,q}=\left(\sum_{i=1}^{n}i^{\frac{q}{p}-1}\right)^{\frac{1}{q}}.$
Now, from the theory of the Riemann integral, we know
$\displaystyle\lim_{n\rightarrow\infty}\frac{\|a^{n}\|_{p,\infty}}{\|a^{n}\|_{p,q}}$ $=\displaystyle\lim_{n\rightarrow\infty}\frac{n^{\frac{1}{p}}}{\displaystyle\left(\sum_{i=1}^{n}i^{\frac{q}{p}-1}\right)^{\frac{1}{q}}}$
$=\displaystyle\frac{1}{\displaystyle\left(\lim_{n\rightarrow\infty}\sum_{i=1}^{n}\left(\frac{i}{n}\right)^{\frac{q}{p}-1}\frac{1}{n}\right)^{\frac{1}{q}}}$
$=\displaystyle\left(\int_{0}^{1}x^{\frac{q}{p}-1}dx\right)^{-\frac{1}{q}}$
$=\left(\frac{q}{p}\right)^{\frac{1}{q}}.$
Hence $\{a^{n}\}_{n=1}^{\infty}$ satisfies (3.1) and we proved $\|I\|=\left(\frac{q}{p}\right)^{\frac{1}{q}}$. ∎
4. Examples
We shall now present examples of embeddings between Lorentz sequence spaces which illustrate Theorem  2.3.
Proposition 4.1.
Let $p,q\in(0,\infty]$ and $\min\{p,q\}<\infty$. Then the embedding operator $I\colon\ell^{p,q}\rightarrow c_{0}$ is maximally non-compact.
Proof.
  1. 1.
    $\ell^{p,q}\subset c_{0}$ from Lemma 3.1.
  2. 2.
    We find the norm of operator $I$. Fix $a=\{a_{n}\}_{n=1}^{\infty}\in\ell_{p,q}$.
    1. a)
      Let $q=\infty$. Then $\|I(a)\|=\|a\|_{\infty}=\sup_{n\in\mathbb{N}}|a_{n}|\leq\sup_{n\in\mathbb{N}}\{n^{1/p}a_{n}^{*}\}=\|a\|_{p,q}$, where the inequality results from the fact that $n^{1/p}\geq 1$ for all $n\in\mathbb{N}$ and the Definition 1.6 of the decreasing rearrangement of $a$.
    2. b)
      Let $q<\infty$. Then $\|I(a)\|=\|a\|_{\infty}=\sup_{n\in\mathbb{N}}|a_{n}|\leq\left(\displaystyle\sum_{n=1}^{\infty}(a_{n}^{*})^{q}n^{\frac{q}{p}-1}\right)^{\frac{1}{q}}=\|a\|_{p,q}$, where the inequality results from the Definition 1.6 of the decreasing rearrangement of $a$ and because $\sup_{n\in\mathbb{N}}|a_{n}|=a_{1}^{*}=1^{\frac{q}{p}-1}(a_{1}^{*})^{q}$.
    Thus, $\|I\|\leq 1$ for any $p\in[1,\infty)$, $q\in[1,\infty]$. Now we will show that $I$ achieves $1$ on the closed unit ball. For an arbitrary $e^{j}$ we have $(e^{j})^{*}=e^{1}$, therefore $\|e^{j}\|_{p,q}=\|e^{1}\|_{p,q}=1$, thus $e^{j}\in B_{\ell_{p,q}}$ and also $\|I(e^{j})\|_{\infty}=\|e^{j}\|_{\infty}=1$, hence $\|I|\|=1$.
So, from Theorem 2.3 the embedding operator $I$ is maximally non-compact and $\alpha(I)=\|I\|=1$. ∎
Proposition 4.2.
Let $0<p_{1}<p_{2}\leq\infty$, $q_{1},q_{2}\in[1,\infty]$, $\min\{p_{2},q_{2}\}<\infty$ and $\ell^{p_{1},q_{1}}$, $\ell^{p_{2},q_{2}}$ Lorentz sequence spaces. Then the embedding operator $I\colon\ell^{p_{1},q_{1}}\rightarrow\ell^{p_{2},q_{2}}$ is maximally non-compact.
Proof.
We have
  1. 1.
    $\ell^{p_{1},q_{1}}\hookrightarrow\ell^{p_{2},q_{2}}$ and $0<\|I\|<\infty$ from Theorem 3.4,
  2. 2.
    $\ell^{p_{1},q_{1}}$, $\ell^{p_{2},q_{2}}\subset c_{0}$ from Lemma 3.1.
So, the prerequisites of Theorem 2.3 are satisfied, the embedding operator $I$ is maximally non-compact and $\alpha(I)=\|I\|$, where the values of $\|I\|$ for some individual relations between $p_{1}$, $q_{1}$ and $q_{2}$ are computed in Theorem 3.5 or Theorem 3.6. ∎
We will now present examples of embeddings of Lorentz sequence spaces into $\ell^{\infty}$ which are not maximally non-compact. This will be shown by finding their span $\sigma_{\ell^{p,q}}$ and applying Theorem 2.4.
Proposition 4.3.
Let $p\in(0,\infty]$, $q\in(0,\infty]$. Then the embedding operator $I\colon\ell^{p,q}\rightarrow\ell^{\infty}$, $\|I\|=1$ is not maximally non-compact. More precisely:
  1. 1.
    $\alpha(I)\leq 2^{-1/q}<1$ when $1\leq q<\infty$ and $p\leq q$,
  2. 2.
    $\alpha(I)\leq 2^{-1}<1$ when $0<q<1$ and $p\leq q$,
  3. 3.
    $\alpha(I)\leq 2^{-1}(1+2^{-1/p})<1$ when $q=\infty$ and $p<q$,
  4. 4.
    $\alpha(I)\leq 2^{-1}(1+2^{-1/p})<1$ when $q<p$.
Proof.
We have $\ell^{p,q}\subset c_{0}\subset\ell^{\infty}$ from Lemma 3.1 and I=1||I||=1 follows from the computations in Example 4.1
  1. 1.
    Let $1\leq q<\infty$ and $p\leq q$. Denote $\sigma=2^{1-1/q}$. Then $\|I\|\leq\sigma<2\|I\|$. Assume $y\in B_{\ell^{p,q}}$, then $y\in B_{\ell^{\infty}}$ since $\|I\|=1$, hence $|y_{j}|\leq 1\leq\sigma$. We claim that
    (4.1) $\sup y-\inf y\leq\sigma.$
    Indeed, given $\varepsilon>0$, we find from the definition of supremum and infimum $s,i\in\mathbb{N}$ such that $y_{s}>\sup y-\varepsilon$ and $y_{i}<\inf y+\varepsilon$. We get
    $1$ $\geq\|y\|_{p,q}$
    $=\left(\displaystyle\sum_{n=1}^{\infty}(y_{n}^{*})^{q}n^{\frac{q}{p}-1}\right)^{\frac{1}{q}}$
    $\geq\left(\displaystyle\sum_{n=1}^{\infty}(y_{n}^{*})^{q}\right)^{\frac{1}{q}}=\left(\displaystyle\sum_{n=1}^{\infty}|y_{n}|^{q}\right)^{\frac{1}{q}}$
    $\geq(|y_{s}|^{q}+|y_{i}|^{q})^{1/q}$
    $\geq 2^{1/q-1}(|y_{s}|+|y_{i}|)$
    $>\frac{1}{\sigma}(\sup y-\inf y-2\varepsilon),$
    where the second inequality follows from our prerequisite $p\leq q$ thus $n^{\frac{q}{p}-1}\geq 1$ and the fourth follows from the Lemma 1.14
    Now when sending $\varepsilon\rightarrow 0_{+}$ we get our claim. Therefore σ_ℓ_p,q ≤σ¡ 2∥I∥ and all the prerequisites of Theorem 2.4 are satisfied, so $I$ is not maximally non-compact and $\alpha(I)\leq\sigma/2<1=\|I\|$.
  2. 2.
    Let $0<q<1$ and $p\leq q$. Denote $\sigma=1$. Then $\|I\|\leq\sigma<2\|I\|$. Analogously to the previous case we assume $y\in B_{\ell^{p,q}}$, hence $|y_{j}|\leq 1\leq\sigma$ and we claim that
    $\sup y-\inf y\leq 1.$
    Indeed, given $\varepsilon>0$, we find $s,i\in\mathbb{N}$ such that $y_{s}>\sup y-\varepsilon$ and $y_{i}<\inf y+\varepsilon$. We get
    $1$ $\geq\|y\|_{p,q}$
    $=\left(\displaystyle\sum_{n=1}^{\infty}(y_{n}^{*})^{q}n^{\frac{q}{p}-1}\right)^{\frac{1}{q}}$
    $\geq\left(\displaystyle\sum_{n=1}^{\infty}(y_{n}^{*})^{q}\right)^{\frac{1}{q}}=\left(\displaystyle\sum_{n=1}^{\infty}|y_{n}|^{q}\right)^{\frac{1}{q}}$
    $\geq(|y_{s}|^{q}+|y_{i}|^{q})^{1/q}$
    $\geq|y_{s}|+|y_{i}|$
    $>\sup y-\inf y-2\varepsilon,$
    where the second inequality follows from our prerequisite $p\leq q$ thus $n^{\frac{q}{p}-1}\geq 1$. and the fourth follows from the Lemma 1.15
    Now when sending $\varepsilon\rightarrow 0_{+}$ we get our claim. Therefore σ_ℓ_p,q ≤1 ¡ 2∥I∥ and all the prerequisites of Theorem 2.4 are satisfied, so $I$ is not maximally non-compact and $\alpha(I)\leq 1/2<1=\|I\|$.
  3. 3.
    Now let $q=\infty$. We will proceed with the proof similarly to the way we did in the first part. The only difference is that we set $\sigma=1+2^{-1/p}$. Assume $y\in B_{\ell^{p,\infty}}$. We again need to verify that the inequality
    (4.2) $\sup y-\inf y\leq\sigma.$
    Let $y\in B_{\ell^{p,\infty}}$, then
    (4.3) $y||_{p,\infty}=\displaystyle\sup_{n\in\mathbb{N}}\{n^{1/p}y_{n}^{*}\}\leq 1\text{, hence }\displaystyle\sup_{n\in\mathbb{N}}\{y_{n}\}\leq 1.$
    Inequality (4.3) implies that the decreasing rearrangement $y^{*}$ of $y$ must satisfy $y_{n}^{*}\leq n^{-1/p}$ for all $n\in\mathbb{N}$. So every sequence from $B_{\ell^{p,\infty}}$ has its decreasing rearrangement bounded from above by the decreasing sequence $\{n^{-1/p}\}_{n=1}^{\infty}$ and from below by $\{-n^{-1/p}\}_{n=1}^{\infty}$. We shall now distinguish two cases.
    1. a)
      If $\sup y\leq 2^{-1/p}$, then from the observation above $\inf y\geq-1$ and inequality (4.2) holds.
    2. b)
      If $\sup y>2^{-1/p}$, then from the observations above $y_{n}^{*}$ is bounded by the decreasing sequence $n^{-1/p}$, so there can be only one index $j\in\mathbb{N}$ such that $y_{j}>2^{-1/p}$, therefore $y_{1}^{*}=y_{j}$ and we get that for all $i\neq j\colon|y_{i}|\leq 2^{-1/p}$, which implies that $\inf y\geq-2^{-1/p}$. Thus $\sup y-\inf y\leq 1-(-2^{-1/p})=\sigma$ holds for all $y\in B_{\ell^{p,q}}$, and inequality (4.2) follows.
    We computed that $\sigma_{\ell_{p,q}}\leq\sigma=1+2^{-1/p}<2\|I\|$. Now again from Theorem 2.4 we get that $I$ is not maximally non-compact and $\alpha(I)\leq\varrho=2^{-1}(1+2^{-1/p})<1$.
  4. 4.
    Now let $q<p$. By the second part of Theorem 3.3 $\ell_{p,q}\hookrightarrow\ell_{p,\infty}$ and $\|y\|_{p,\infty}\leq\|y\|_{p,q}$ for all $y\in\ell_{p,q}$. Now assume $y\in B_{\ell^{p,q}}$. Then $y\in B_{\ell^{p,\infty}}$, so from the proof of the previous case, where $q=\infty$, we again get $y_{n}^{*}\leq n^{-1/p}$ and $\sigma_{\ell_{p,q}}\leq 1+2^{-1/p}<2\|I\|$. Now again from Theorem 2.4 we get that $I$ is not maximally non-compact and α(I) ≤ϱ= 2^-1(1+2^-1/p) ¡ 1.
Acknowledgment. This research was in part supported by the grant no. 23-04720S of the Czech Science Foundation.